Write in terms of sine and cosine:
cos^2 θ(tan^2 θ+1) =
(secθ-cosθ)/(sinθ) =
-----------------------------------------------------
Simplify the trig expression to a power of a single trig function:
(sec^2x-1)/(sec^2x)
Please be descriptive on how these are solved
Thank you

Answer by stanbon(75887)

Write in terms of sine and cosine:
cos^2 θ(tan^2 θ+1) = cos^2*sec^2 = 1
---------------------
(secθ-cosθ)/(sinθ) = ((/cos)-cos)/sin = ((1-cos^2)/cos) / sin
---
= (sin^2/cos)/sin = sin/cos = tan(theta)
------------------------------------------------
-----------------------------------------------------
Simplify the trig expression to a power of a single trig function:
(sec^2x-1)/(sec^2x) = (1 - (1/sec^2(x)) = (1-cos^2(x)) = sin^2(x)
==================================
Please be descriptive on how these are solved
Cheers,
Stan H.
----------